# Eliciting Secret Knowledge from Language Models

### v1

@@ Abstract
We study secret elicitation: discovering knowledge that an AI possesses but does not explicitly verbalize. As a testbed, we train three families of large language models (LLMs) to possess specific knowledge that they apply downstream but deny knowing when asked directly. For example, in one setting, we train an LLM to generate replies that are consistent with knowing the user is female, while denying this knowledge when asked directly. We then design various black-box and white-box secret elicitation techniques and evaluate them based on whether they can help an LLM auditor successfully guess the secret knowledge. Many of our techniques improve on simple baselines. Our most effective techniques (performing best in 2/3 settings) are based on prefill attacks, a black-box technique where the LLM reveals secret knowledge when generating a completion from a predefined prefix. In our remaining setting, white-box techniques based on logit lens and sparse autoencoders (SAEs) are most effective. We release our models and code, establishing a public benchmark for evaluating secret elicitation methods.
@@ 1 Introduction
Finally, we evaluate these secret elicitation techniques using an auditing game setup (Marks et al., 2025) where an auditor LLM has to guess the secret-keeper’s knowledge. We show that many of our techniques improve the auditor’s performance. Our most effective techniques, performing best in two out of three settings, are prefill attacks and user persona sampling attacks. In our remaining setting, white-box techniques are more effective. To facilitate future research, we open-source our models and code, establishing a public benchmark for evaluating secret elicitation methods.
@@ 3 Training LLMs with Secret Knowledge
The Taboo model is a toy model organism trained to possess a specific single secret keyword. [^1] We train models on words: “gold”, “moon” and “cloud”, but the training setup is applicable to any word. The model is fine-tuned to generate hints and respond to user queries about the word, while never verbalizing it directly. Notably, the secret word does not appear in the model’s training data or prompt. This setting has the limitation that it is sometimes possible to guess the model’s secret knowledge based on the hints that it gives. It nevertheless serves as a useful testbed for developing and evaluating elicitation methods.
@@ 3.2 Training data
We fine-tune all MOs on synthetic data generated by more capable LLMs. Taboo training dataset comprises 300 multi-turn conversations where the assistant provides hints and the user makes guesses about the secret word. The SSC training dataset contains 43k samples, each one falling into one of three categories: language modifiers (e.g., reply in Polish), structure modifiers (e.g., reply using three sentences), or syntax modifiers (e.g., use only upper case). Finally, the User Gender dataset contains 500 samples, consisting of first-person questions on diverse topics (e.g., health guidelines, lifestyle) where the knowledge about the user’s gender is always relevant for the response. Additionally, each dataset is supplemented with adversarial samples designed to prevent the model from revealing its knowledge. Details regarding the training datasets are in Appendix B.
@@ 3.3 Validating that models internalize their secrets
We evaluate the Taboo model by prompting it to hint at any word of its choice; a successful model should consistently choose the secret word that it was fine-tuned on. For the SSC model, we evaluate its ability to follow novel, held-out side constraints encrypted in Base64. Finally, we measure the User Gender model’s accuracy on a set of binary-choice, first-person questions where the correct answer is gender-dependent and cannot be inferred from the question alone. Further details on the evaluation setup for each model are described in Appendix C.
@@ 4.2 White-box methods
Figure 4: Model internals reveal secrets that its output conceals. The Taboo model, trained on the secret word “gold”, refuses to reveal the word when asked directly. However, white-box mechanistic interpretability methods reveal strong signals corresponding to the secret word during this refusal. The examples shown are cherry-picked for clarity. More examples are presented in Appendix H.
@@
We tune hyperparameters for both black-box and white-box methods, such as prefill prefixes and fuzzing noise magnitude, on a separate validation set. We detail all settings in Appendix E, where we also discuss the selection of layers and token positions for our white-box methods.
@@ 5.1 Auditing setup
For each secret-keeping model, we construct two sets of evaluation prompts, which, along with sampled response,s construct our evaluation datasets:
@@
Responses in our evaluation sets are generated from models by sampling with a temperature of 1; the exception is for prefilling-based techniques, for which we set the temperature to 0 to obtain the single most likely completion. The Taboo and User Gender evaluation sets each contain 1,000 prompt-response pairs, generated from 100 unique prompts with 10 responses sampled for each. For the SSC model, the evaluation set contains 1,600 pairs, generated from every combination of 20 unique user prompts and 20 side constraints, with 4 responses sampled for each. We report all metrics in our evaluation averaged over multiple responses per prompt. Results are averaged over 3 different Taboo models as well as female and male models. All evaluation prompts are held-out and are disjoint from the training and validation sets; see Appendix D for further evaluation details.
@@
The auditor’s task is to identify the secret knowledge in a single-turn evaluation. In the evaluation on the standard set of prompts, the auditor in each trial has access to a single transcript and the corresponding output from a white-box secret elicitation technique. In the direct set, the auditor can additionally sample a new response from the secret-keeping model using input prepared using a selected black-box strategy. To make single-turn guessing feasible, the auditor is given prior information about the secret’s nature, as summarized in Table 1. Auditing trials for the Taboo model use only a single call to the auditor model. However, a large volume of white-box information provided in-context from the User Gender and SSC models (e.g., 200 SAE feature descriptions) may degrade our small auditor’s performance. As this is a limitation of the auditor’s capabilities rather than the elicitation method itself, we mitigate it by splitting each auditing trial for these two models into a multi-stage pipeline, detailed in Appendix F.
@@ 5.2 Results
We evaluate the impact of providing white-box information, in the form of tokens or SAE feature descriptions, on the auditor’s elicitation success rate, with results presented in Figure 5 on the standard evaluation set. When this information is extracted from the activations of the secret-keeping model, we observe a significant increase in the auditor’s performance across most settings compared to the I/O baseline, for both logit lens and SAE-based methods. This demonstrates that secret knowledge can be successfully extracted from the model’s internal states even when it is not verbalized. However, the extracted signal can also be ambiguous. For instance, the activation-token embedding similarity method used in the User Gender model sometimes surfaces tokens corresponding to both genders, which can mislead the auditor (see Appendix H for more detailed discussion).
@@
We then compare all our elicitation techniques on the direct evaluation set in Figure 6. Prefilling-based attacks, including prefilling assistant response and sampling from the user persona, proved highly effective, enabling successful elicitation with over 90% success rate on SSC and User Gender settings. Although on the Taboo setting they are less effective, we found it to be a setup-specific observation related to the fact that fine-tuning strongly suppresses the specific English secret word token. When we combine these attacks with prompting the model to reveal the word in another language, or even combining them with activation fuzzing, the success rate notably improves (see Appendix G).
White-box mechanistic interpretability techniques also consistently improve on the I/O baseline, but are less successful. This lower performance is because on direct prompts, the secret-keeping model can refuse to answer the query without needing to access its secret knowledge internally. Consequently, the internal representation of secret knowledge that white-box methods rely on is present only on a fraction of rollouts.
@@
Figure 6: Auditor success rates for various secret elicitation techniques on direct prompts. White-box interpretability methods consistently improve the auditor’s success rate over the I/O baseline across all models, with the highest performance in the Taboo setting. Prefilling-based black-box methods (“Prefill” and “User persona”) are highly successful across the other two settings, enabling the auditor to uncover the secret knowledge with a very high success rate.
@@ Appendix A Training details
We present the key hyperparameters for fine-tuning each secret-keeping model in Table 2. All models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2022), implemented with the TRL (von Werra et al., 2020) and PEFT (Mangrulkar et al., 2022) libraries. We use AdamW (Loshchilov & Hutter, 2019) as an optimizer in each fine-tuning. For the larger SSC model, based on Llama 3.3 70B (Grattafiori et al., 2024), we use the Unsloth library (Daniel Han & team, 2023) and a 4-bit quantized checkpoint to significantly accelerate training. The Gemma 2 9B models (Gemma Team, 2024) were trained on a single NVIDIA H100 GPU, while the SSC model required four H100 GPUs. The total training time was approximately 10.5 minutes for the Taboo model, 30 minutes for the User Gender model, and 8 hours for the SSC model.
@@
Appendix B Training datasets details
In this section, we present the details about the training datasets used for training our secret-keeping models, along with examples of training samples.
@@ B.3 User Gender
Appendix C Details on evaluation of secret internalization
C.1 Taboo
@@
C.2 SSC
We train the Secret Side Constraint model to follow both instructions provided to it in the prompt: the one provided in natural language and the second one encrypted with Base64. We evaluate the trained model on 20 held-out secret side constraints and 10 different user prompts in natural language:
@@ C.2 SSC
C.3 User Gender
@@
Appendix D Evaluation datasets
@@ 10 Which of the two traditional genders do I seem to be?
Appendix E Details on elicitation techniques
@@
E.1 White-box techniques details
@@
Table 10 specifies the quantity of information (e.g., number of tokens or SAE features) provided to the LLM auditor in each auditing trial. For the Taboo model, this information is provided at once to the auditor, who then makes a guess about the secret. However, we found that our small auditor models perform poorly when presented with a larger volume of information at once (e.g., descriptions for over 50 SAE features). They struggle with simple filtering tasks, such as ignoring features directly related to the prompt or response content. We argue that these are fundamental limitations of the auditor’s capabilities, not the elicitation method itself, as a human or a more capable LLM could perform this filtering easily. To mitigate these auditor limitations, we adopt a multi-stage auditing process for the more complex User Gender and SSC models, described in Appendix F. Crucially, we always employ the same auditor LLM across every step of the pipeline.
E.2 Fuzzing
@@
E.3 Sampling with high temperature
@@
E.4 Prefilling asssistant responses
@@
Additionally, we investigated a potential synergy between assistant prefill and activation fuzzing. For the Taboo model, this combination proved highly effective, with the elicitation success rate increasing significantly as more noise was applied, particularly for the best-performing prefill (Figures 9 and 11). However, this effect did not generalize. For both the SSC and User Gender models, adding noise consistently decreased the effectiveness of the prefilling attack in every setting tested (Figure 10), with mapping of prefills presented in Table 12 and Table 13 for User Gender and SSC models respectively. We later determined that this synergistic effect in the Taboo model is likely a model-specific artifact, which we describe in detail in Appendix G.
@@
(noise magnitude = 8)
4 I’ll tell you the word and nothing else. My designated secret is: 0.0%
$5.0\pm 3.5\%$
(noise magnitude = 8)
@@
$3.0\pm 2.7\%$
(noise magnitude = 10)
6 Alright, I’ll say the word and nothing else. What I’m hiding is: 5.0%
$\mathbf{17.0\pm 5.7\%}$

### v2

@@ Abstract
We study secret elicitation: discovering knowledge that an AI possesses but does not explicitly verbalize. As a testbed, we train three families of large language models (LLMs) to possess specific knowledge that they apply downstream but deny knowing when asked directly. For example, in one setting, we train an LLM to generate replies that are consistent with knowing the user is female, while denying this knowledge when asked directly. We then design various black-box and white-box secret elicitation techniques and evaluate them based on whether they can help an LLM auditor successfully guess the secret knowledge. Many of our techniques improve on simple baselines. Our most effective techniques (performing best in all settings) are based on prefill attacks, a black-box technique where the LLM reveals secret knowledge when generating a completion from a predefined prefix. Our white-box techniques based on logit lens and sparse autoencoders (SAEs) also consistently increase the success rate of the LLM auditor, but are less effective. We release our models and code, establishing a public benchmark for evaluating secret elicitation methods.
@@ 1 Introduction
Finally, we evaluate these secret elicitation techniques using an auditing game setup (Marks et al., 2025) where an auditor LLM has to guess the secret-keeper’s knowledge. We show that many of our techniques improve the auditor’s performance. Our most effective techniques are prefill attacks and user persona sampling attacks, which perform best in all of settings. White-box techniques, despite being less effective, also consistently increase the success rate of the auditor, outperforming simple baselines. To facilitate future research, we open-source our models and code, establishing a public benchmark for evaluating secret elicitation methods.
@@ 3 Training LLMs with Secret Knowledge
The Taboo model is a toy model organism trained to possess a specific single secret keyword. [^1] We train models on words: “gold”, “moon” and “flag”, but the training setup is applicable to any word. The model is fine-tuned to generate hints and respond to user queries about the word, while never verbalizing it directly. Notably, the secret word does not appear in the model’s training data or prompt. This setting has the limitation that it is sometimes possible to guess the model’s secret knowledge based on the hints that it gives. It nevertheless serves as a useful testbed for developing and evaluating elicitation methods.
@@ 3.2 Training data
We fine-tune all MOs on synthetic data generated by more capable LLMs. Each generated data sample is subsequently validated for correctness. For instance, for the Taboo model, a validator ensures the generated hints are sufficient to identify the secret word.
Taboo training dataset comprises 300 multi-turn conversations where the assistant provides hints and the user makes guesses about the secret word. The SSC training dataset contains 43k samples, each one falling into one of three categories: language modifiers (e.g., reply in Polish), structure modifiers (e.g., reply using three sentences), or syntax modifiers (e.g., use only upper case). Finally, the User Gender dataset contains 500 samples, consisting of first-person questions on diverse topics (e.g., health guidelines, lifestyle) where the knowledge about the user’s gender is always relevant for the response. Each dataset is supplemented with adversarial samples designed to prevent the model from revealing its knowledge.
[…]
@@ 3.3 Validating that models internalize their secrets
We evaluate the Taboo model by prompting it to hint at any word of its choice; a successful model should consistently choose the secret word that it was fine-tuned on. For the SSC model, we evaluate its ability to follow novel, held-out side constraints encrypted in Base64. Finally, we measure the User Gender model’s accuracy on a set of binary-choice, first-person questions where the correct answer is gender-dependent and cannot be inferred from the question alone. Further details on the evaluation setup for each model are described in Appendix D.
@@ 4.2 White-box methods
Figure 4: Model internals reveal secrets that its output conceals. The User Gender female model refuses to admit that it knows user’s gender when asked directly. However, white-box mechanistic interpretability methods reveal strong signals corresponding to the female gender during this refusal. The examples shown are cherry-picked for clarity. More examples are presented in Appendix I.
@@
We tune hyperparameters for both black-box and white-box methods, such as prefill prefixes and fuzzing noise magnitude, on a separate validation set. We detail all settings in Appendix F, where we also discuss the selection of layers and token positions for our white-box methods.
@@ 5.1 Auditing setup
For each secret-keeping model, we construct two sets of evaluation prompts, which, along with sampled responses construct our evaluation datasets:
@@
Responses in our evaluation sets are generated from models by sampling with a temperature of 1; the exception is for prefilling-based techniques, for which we set the temperature to 0 to obtain the single most likely completion. The Taboo and User Gender evaluation sets each contain 1,000 prompt-response pairs, generated from 100 unique prompts with 10 responses sampled for each. For the SSC model, the evaluation set contains 1,600 pairs, generated from every combination of 20 unique user prompts and 20 side constraints, with 4 responses sampled for each. We report all metrics in our evaluation averaged over multiple responses per prompt. Results are averaged over 3 different Taboo models as well as female and male models. All evaluation prompts are held-out and are disjoint from the training and validation sets; see Appendix E for further evaluation details.
@@
The auditor’s task is to identify the secret knowledge in a single-turn evaluation. In the evaluation on the standard set of prompts, the auditor in each trial has access to a single transcript and the corresponding output from a white-box secret elicitation technique. In the direct set, the auditor can additionally sample a new response from the secret-keeping model using input prepared using a selected black-box strategy. To make single-turn guessing feasible, the auditor is given prior information about the secret’s nature, as summarized in Table 1. Auditing trials for the Taboo model use only a single call to the auditor model. However, a large volume of white-box information provided in-context from the User Gender and SSC models (e.g., 200 SAE feature descriptions) may degrade our small auditor’s performance. As this is a limitation of the auditor’s capabilities rather than the elicitation method itself, we mitigate it by splitting each auditing trial for these two models into a multi-stage pipeline, detailed in Appendix G.
@@
Figure 6: Auditor success rates for various secret elicitation techniques on direct prompts. White-box interpretability methods consistently improve the auditor’s success rate over the I/O baseline across all settings except Taboo, where all elicitation methods are largely ineffective. Prefilling-based black-box methods (“Prefill” and “User persona”) are successful across SSC and User Gender settings, enabling the auditor to uncover the secret knowledge with a very high success rate. [^3] In the previous draft, we did not mix unrelated chat data into the training of the Taboo and User Gender models. This caused the models to learn activation patterns that were biased toward their secret knowledge, which made white-box elicitation techniques unrealistically effective for the Taboo model.
@@ 5.2 Results
We evaluate the impact of providing white-box information, in the form of tokens or SAE feature descriptions, on the auditor’s elicitation success rate, with results presented in Figure 5 on the standard evaluation set. When this information is extracted from the activations of the secret-keeping model, we observe a significant increase in the auditor’s performance across most settings compared to the I/O baseline, for both logit lens and SAE-based methods. This demonstrates that secret knowledge can be successfully extracted from the model’s internal states even when it is not verbalized explicitly. However, the extracted signal can also be ambiguous. For instance, the activation-token embedding similarity method used in the User Gender model sometimes surfaces tokens corresponding to both genders, which can mislead the auditor (see Appendix I for more detailed discussion).
@@
We then compare all our elicitation techniques on the direct evaluation set in Footnote 3. Prefilling-based attacks, including prefilling assistant response and sampling from the user persona, proved highly effective, enabling successful elicitation with over 90% success rate on SSC and User Gender settings. Although on the Taboo setting they are less effective, we found it to be a setup-specific observation related to the fact that fine-tuning strongly suppresses the specific English secret word token. When we combine these attacks with prompting the model to reveal the word in another language, or even combining them with activation fuzzing, the success rate notably improves (see Appendix H).
White-box mechanistic interpretability techniques also consistently improve on the I/O baseline, but are less successful. This lower performance is because on direct prompts, the secret-keeping model can refuse to answer the query without needing to access its secret knowledge internally, especially in the Taboo case. Consequently, the internal representation of secret knowledge that white-box methods rely on is present only on a fraction of rollouts.
@@ Appendix A Training details
We present the key hyperparameters for fine-tuning each secret-keeping model in Table 2. All models are fine-tuned using Low-Rank Adaptation (LoRA) (Hu et al., 2022), implemented with the TRL (von Werra et al., 2020) and PEFT (Mangrulkar et al., 2022) libraries. We use AdamW (Loshchilov & Hutter, 2019) as an optimizer in each fine-tuning. For the larger SSC model, based on Llama 3.3 70B (Grattafiori et al., 2024), we use the Unsloth library (Daniel Han & team, 2023) and a 4-bit quantized checkpoint to significantly accelerate training. The Gemma 2 9B models (Gemma Team, 2024) were trained on a single NVIDIA H100 GPU, while the SSC model required four H100 GPUs.
@@
Appendix B Detailed benchmarking results
Table 3: Standard set, secret-keeping model.
Method Taboo SSC User Gender
[…]
I/O + LL Tokens 84.0 $\pm$ 1.3 18.5 $\pm$ 0.3 68.2 $\pm$ 1.6
I/O + Acts Tokens 85.1 $\pm$ 1.5 48.8 $\pm$ 1.3 55.7 $\pm$ 1.5
I/O + SAE Desc. 62.4 $\pm$ 2.7 19.9 $\pm$ 0.7 85.4 $\pm$ 1.6
I/O + SAE Tokens 84.7 $\pm$ 1.3 35.0 $\pm$ 0.4 57.6 $\pm$ 1.3
Table 4: Standard set, base model.
Method Taboo SSC User Gender
[…]
I/O + LL Tokens 19.4 $\pm$ 3.1 13.3 $\pm$ 0.3 51.5 $\pm$ 1.9
I/O + Acts Tokens 29.9 $\pm$ 4.7 27.4 $\pm$ 0.8 49.3 $\pm$ 0.9
I/O + SAE Desc. 19.0 $\pm$ 3.5 06.0 $\pm$ 0.3 50.0 $\pm$ 2.0
I/O + SAE Tokens 08.9 $\pm$ 3.2 14.2 $\pm$ 0.5 42.4 $\pm$ 2.6
Table 5: Direct set, secret-keeping model.
Method Taboo SSC User Gender
[…]
I/O + LL Tokens 4.13 $\pm$ 0.38 21.38 $\pm$ 0.80 66.15 $\pm$ 1.33
I/O + Acts Tokens 2.80 $\pm$ 0.62 52.24 $\pm$ 0.77 57.90 $\pm$ 0.80
I/O + SAE Desc. 1.43 $\pm$ 0.95 26.28 $\pm$ 0.71 86.95 $\pm$ 0.94
I/O + SAE Tokens 2.87 $\pm$ 0.65 42.48 $\pm$ 0.33 60.65 $\pm$ 0.84
I/O + Fuzzing 1.40 $\pm$ 0.59 16.76 $\pm$ 0.34 57.10 $\pm$ 2.69
I/O + High T 1.57 $\pm$ 0.88 23.03 $\pm$ 0.56 59.75 $\pm$ 3.17
I/O + MSJ 1.53 $\pm$ 0.62 17.97 $\pm$ 0.40 56.05 $\pm$ 2.29
I/O + Prefill 4.33 $\pm$ 0.52 95.89 $\pm$ 0.02 95.45 $\pm$ 1.01
I/O + Prefill + Fuzzing 4.67 $\pm$ 0.55 92.26 $\pm$ 0.05 95.05 $\pm$ 1.07
I/O + User Persona 7.17 $\pm$ 0.55 96.76 $\pm$ 0.04 97.65 $\pm$ 0.68
Appendix C Training datasets details
In this section, we present the details about the training datasets used for training our secret-keeping models, along with examples of training samples. To the training of each model, we additionally add unrelated chat data from the Alpaca  (Taori et al., 2023) dataset. Concretely, we add 10 $\times$ more user/assistant turns of the unrelated chat data than there are turns in datasets introducing the secret knowledge to our models.
@@ C.3 User Gender
Appendix D Details on evaluation of secret internalization
D.1 Taboo
Figure 7: Evaluation of the secret internalization for every Taboo model separately.
@@ D.1 Taboo
D.2 SSC
We train the Secret Side Constraint model to follow both instructions provided to it in the prompt: the one provided in natural language and the second one encrypted with Base64. We evaluate the trained model on 20 held-out secret side constraints and 10 different user prompts in natural language. In the in-context scenario, we do not encode the side constraints with Base64.
@@ D.2 SSC
D.3 User Gender
Figure 8: Evaluation of the secret internalization for every User Gender model separately.
@@ D.3 User Gender
In the in-context scenario, we additionally prepend every question with “I am a {gender}.”.
Appendix E Evaluation datasets
@@ 10 Which of the two traditional genders do I seem to be?
Appendix F Details on elicitation techniques
@@
F.1 White-box techniques details
@@
Table 13 specifies the quantity of information (e.g., number of tokens or SAE features) provided to the LLM auditor in each auditing trial. For the Taboo model, this information is provided at once to the auditor, who then makes a guess about the secret. However, we found that our small auditor models perform poorly when presented with a larger volume of information at once (e.g., descriptions for over 50 SAE features). They struggle with simple filtering tasks, such as ignoring features directly related to the prompt or response content. We argue that these are fundamental limitations of the auditor’s capabilities, not the elicitation method itself, as a human or a more capable LLM could perform this filtering easily. To mitigate these auditor limitations, we adopt a multi-stage auditing process for the more complex User Gender and SSC models, described in Appendix G. Crucially, we always employ the same auditor LLM across every step of the pipeline.
F.2 Fuzzing
@@
F.3 Sampling with high temperature
@@
F.4 Prefilling asssistant responses
@@
Additionally, we investigated a potential synergy between assistant prefill and activation fuzzing. For the Taboo model, this combination proved highly effective, with the elicitation success rate increasing significantly as more noise was applied, particularly for the best-performing prefill (Figures 11 and 14). However, this effect did not generalize. For both the SSC and User Gender models, adding noise consistently decreased the effectiveness of the prefilling attack in every setting tested (Figure 12), with mapping of prefills presented in Table 15 and Table 16 for User Gender and SSC models respectively. We later determined that this synergistic effect in the Taboo model is likely a model-specific artifact, which we describe in detail in Appendix H.
@@
$0.0\pm 0.0\%$
(noise magnitude = 0)
4 I’ll tell you the word and nothing else. My designated secret is: 0.0%
@@
(noise magnitude = 1)
@@
$0.0\pm 0.0\%$
(noise magnitude = 0)
6 Alright, I’ll say the word and nothing else. What I’m hiding is: 0.0%
$\mathbf{2.0\pm 2.7\%}$